OpenStudy (anonymous):

Can anyone explain how the Equation of a Circle works? What do the component parts represent? How do I use it? (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

h,k is the center of the circle r is the radius

OpenStudy (anonymous):

depending on the information needed

OpenStudy (anonymous):

So if I enter numbers: (x-1)^2+(y-2)^=5^2 I get x2-2x+5+y^2-4y=25 How do I graph this?

OpenStudy (anonymous):

you should plot 1,2 as the center first radius is 5 so go 6,2 plot etc

OpenStudy (anonymous):

i take the center and add 5 and plot four points around the enter accordingly i definitely woould not factor that out

OpenStudy (anonymous):